Water Company purchased a new piece of equipment costing $100,000 in 2018. It estimates it will have a residual value of $40,000 and a useful life of 3 years. The machine was expected to produce 60,000 units over its useful life, produced 30,000 units in 2018, and produced 15,000 units in 2019. What amount of depreciation expense should be recorded in 2019 using the units of production method of accounting for depreciation? Enter your answer as a number - no dollar signs and no commas!

alrighty so please only answer this if you are 100% sure you are correct because lately i have not been getting correct answers and i don't pay $10 a month for other people to do my homework incorrectly for me.
Answer #1

Depreciation expense per unit = (Original cost-Salvage value)/Estimated production

= (100000-40000)/60000

Depreciation expense per unit = 1 per unit

Depreciation expense for 2019 = 15000*1 = $15000